OpenStudy (anonymous):

Over two summers, Ray saved $900 and $500. The polynomial 900x2 + 500x represents his savings at the beginning of the third year, where x is the growth factor. (The interest rate r is x - 1.) What is the interest rate he needs to start the third year with $1600 in savings?

OpenStudy (anonymous):

@rajee_sam @dan815 @Numb3r1

OpenStudy (anonymous):

That means 900x^2+500x=1600, which means 900x^2+500x-1600=0. Factor and solve!

OpenStudy (anonymous):

I should look for x

OpenStudy (rajee_sam):

after you find x the interest rate should be x-1

OpenStudy (anonymous):

okay 1 sec

OpenStudy (anonymous):

I got 2 solutions

OpenStudy (rajee_sam):

one will be negative and one will be +ve

OpenStudy (rajee_sam):

I cannot have negative growth rate so eliminate that

OpenStudy (anonymous):

A. 1.1% B. 0.1% C. 8.4% D. 264%

OpenStudy (anonymous):

these are my solutions http://www.wolframalpha.com/input/?i=900x%5E2%2B500x-1600%3D0

OpenStudy (anonymous):

Just looking at D it is way too high. B is too low. Now, I can look at these and tell which is correct, but it's better to take A and C and see which works with the math.

OpenStudy (anonymous):

That doesn't work.

OpenStudy (anonymous):

That's x^2+7x-16, not 5x.

OpenStudy (rajee_sam):

no 900x^2 + 500x = 1600 so it will be 9x^2 + 5x -16 = 0 not 7x

OpenStudy (rajee_sam):

I did my factoring wrong my bad

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=9x%5E2%2B5x-16%3D0 these are the right solutions

OpenStudy (rajee_sam):

1.1 is you x so r = 1.1 - 1 = .1%

OpenStudy (rajee_sam):

yes your solutions are right

OpenStudy (rajee_sam):

if you look at both your solutions first one is negative. second one is positive. So consider that. if you calculate that it turns out to be 1.1

OpenStudy (rajee_sam):

So Answer is B

OpenStudy (rajee_sam):

if x = 1.1 then r = x-1

OpenStudy (anonymous):

yeah tha't correct I was thinking about that one thanks all

OpenStudy (arihangdu):

Is this correct?